# Qingdao 266237, China

### v1

@@ Abstract
Multi-party private set union (MPSU) protocol enables $m$ $(m>2)$ parties, each holding a set, to collectively compute the union of their sets without revealing any additional information to other parties. There are two main categories of MPSU protocols: The first builds on public-key techniques. All existing works in this category involve a super-linear number of public-key operations, resulting in poor practical efficiency. The second builds on oblivious transfer and symmetric-key techniques. The only existing work in this category is proposed by Liu and Gao (ASIACRYPT 2023), which features the best concrete performance among all existing protocols, despite its super-linear computation and communication. Unfortunately, it does not achieve the standard semi-honest security, as it inherently relies on a non-collusion assumption, which is unlikely to hold in practice. Therefore, the problem of constructing a practical MPSU protocol based on oblivious transfer and symmetric-key techniques in standard semi-honest model remains open. Furthermore, there is no MPSU protocol achieving both linear computation and linear communication complexity, which leaves another unresolved problem. In this work, we resolve these two open problems.
  • We propose the first MPSU protocol based on oblivious transfer and symmetric-key techniques in the standard semi-honest model. This protocol is 4.99.3×4.9-9.3$\times$4.9 - 9.3 × faster than Liu and Gao in the LAN setting. Concretely, our protocol requires only $3.6$ seconds in online phase for 3 parties with sets of $2^{20}$ items each.
@@
We implement our protocols and conduct an extensive experiment to compare the performance of our protocols and the state-of-the-art. To the best of our knowledge, our implementation is the first correct and secure implementation of MPSU that reports on large-size experiments.
@@ 1 Introduction
Over the last decade, there has been growing interest in private set operation (PSO), which consists of private set intersection (PSI), private set union (PSU), and private computing on set intersection (PCSI), etc. Among these functionalities, PSI, especially two-party PSI [PSZ14, KKRT16, PRTY19, CM20, PRTY20, RS21, RR22], has made tremendous progress and become highly practical with extremely fast and cryptographically secure implementations. Meanwhile, multi-party PSI [KMP+17, NTY21, CDG+21, BNOP22] is also well-studied. In contrast, the advancement of PSU has been sluggish until recently, several works proposed efficient two-party PSU protocols [KRTW19, GMR+21, JSZ+22, ZCL+23, CZZ+24a]. However, multi-party PSU has still not been extensively studied. In this work, we focus on PSU in the multi-party setting.
@@
Multi-party private set union (MPSU) enables $m$ $(m>2)$ mutually untrusted parties, each holding a private set of elements, to compute the union of their sets without revealing any additional information. MPSU and its variants have numerous applications, such as information security risk assessment [LV04], IP blacklist and vulnerability data aggregation [HLS+16], joint graph computation [BS05], distributed network monitoring [KS05], building block for private DB supporting full join [KRTW19], private ID [GMR+21] etc.
@@
According to the underlying techniques, existing MPSU protocols can be mainly divided into two categories: The first category, denoted PK-MPSU, is primarily based on public-key techniques, and has been explored in a series of works [KS05, Fri07, VCE22, GNT23]. The drawbacks of these works are that each party has to perform a substantial number of public-key operations, leading to super-linear computation complexity and poor practical efficiency. The second category, denoted SK-MPSU, is primarily based on symmetric-key techniques, and has only one existing work [LG23] to date. This work exhibits much better performance than all prior works. However, it fails to achieve the standard semi-honest security due to its inherent reliance on a non-collusion assumption, assuming the party who obtains the union (we call it leader hereafter) not to collude with other parties. Furthermore, it has super-linear complexity as well. Motivated by the above, we raise the following two questions:
@@
Can we construct a MPSU protocol based on oblivious transfer and symmetric-key operations, without any non-collusion assumptions? Can we construct a MPSU protocol with both linear computation and linear communication complexity [^1] In the context of MPSU, linear complexity means that the complexity for each party scales linearly with the total size of all parties’ sets. In this paper, we consider the balanced setting where each party holds sets of equal size, thus linear complexity means that the complexity for each party scales linearly with both the number of parties $m$ and the set size $n$. Meanwhile, following current conventions in the area of PSO, linear complexity only considers the online phase.?
@@ 1.1 Our Contribution
In this work, we resolve the above two open problems by first presenting a new primitive for MPSU, called batch secret-shared private membership test (batch ssPMT), then employing the batch ssPMT to build a SK-MPSU in the standard semi-honest model, and a PK-MPSU with linear computation and communication complexity. Our contributions are summarized as follows:
1.1.1 Efficient Batch ssPMT.
At the technical core of the state-of-the-art MPSU protocol [LG23] (hereafter referred to as LG protocol) is the multi-query secret-shared private membership test (mq-ssPMT), which dominates both computation and communication costs of LG protocol. In analogy of the relation between batched oblivious pseudorandom function (batch OPRF) [KKRT16] and multi-point oblivious pseudorandom function (multi-point OPRF) [PRTY19, CM20], we abstract a new functionality called batch secret-shared private membership test (batch ssPMT), which allows a fairly efficient construction and can be used to build an alternative to mq-ssPMT in the context of MPSU. Looking ahead, our batch ssPMT serves as a core building block in our two MPSU protocols, and significantly contributes to our speedup compared to LG protocol.
1.1.2 SK-MPSU in Standard Semi-Honest Model.
We generalize random OT (ROT) into multi-party setting, which we call multi-party secret-shared random oblivious transfer (mss-ROT). Based on batch ssPMT and mss-ROT, we propose the first SK-MPSU protocol in the standard semi-honest model. Compared to LG protocol, our SK-MPSU has superior online performance with a 4.99.3×4.9-9.3$\times$4.9 - 9.3 × improvement in the LAN setting.
1.1.3 PK-MPSU with Linear Complexity.
Based on batch ssPMT and multi-key rerandomizable public-key encryption (MKR-PKE) [GNT23], we propose the first MPSU protocol with both linear computation and communication. Our PK-MPSU has the lowest overall communication costs with a factor of 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement compared to LG protocol. It also achieves a 1.85.4×1.8-5.4$\times$1.8 - 5.4 × speedup in terms of overall running time in the WAN setting. Along the way, we find that the PK-MPSU protocol of Gao et al. [GNT23] is insecure against arbitrary collusion and give a practical attack to demonstrate that it necessitates the same non-collusion assumption as LG protocol as well.
@@
1.2 Related Works
We review the existing semi-honest MPSU protocols in the literature.
1.2.1 PK-MPSU.
Kisser and Song [KS05] introduced the first MPSU protocol, based on polynomial representations and additively homomorphic encryption (AHE). This protocol requires a substantial number of AHE operations and high-degree polynomial calculations, so it is completely impractical.
@@ 1.2 Related Works
Vos et al. [VCE22] proposed a MPSU protocol based on the bit-vector representations. The parties collectively compute the union by performing the private OR operations on the bit-vectors, realized by ElGamal encryption. The original version of this protocol is merely applicable for small universes. Even after leveraging divide-and-conquer approach, the protocol requires quadratic computation and communication complexity in the number of parties for the leader and shows poor concrete efficiency reported by [LG23].
Recently, Gao et al. [GNT23] proposed a MPSU protocol based on three newly introduced cryptographic tools, including membership Oblivious Transfer (mOT), conditional Oblivious Pseudorandom Function (cOPRF) and MKR-PKE. This protocol achieves near-linear complexity in the set size and linear complexity in the number of parties, and is the most advanced MPSU in terms of theoretical complexity. Unfortunately, their protocol turns out to be insecure against arbitrary collusion. We propose a practical attack to show that it requires the same non-collusion assumption as LG protocol (see Appendix 0.A for details).
1.2.2 SK-MPSU.
Recently, Liu and Gao [LG23] proposed a practical MPSU protocol based on oblivious transfer and symmetric-key operations. This protocol is several orders of magnitude faster than the prior works. For instance, when computing on datasets of $2^{10}$ element, it is 109×109$\times$109 × faster than [VCE22]. However, their protocol is not secure in the standard semi-honest model.
1.2.3 Other Related Works.
Blanton et al. [BA12] proposed a MPSU protocol based on oblivious sorting and generic MPC. Their work focuses on devising the circuit for MPSU. The heavy dependency on general MPC leads to inefficiency.
@@
Table 1: Asymptotic communication (bits) and computation costs of MPSU protocols in the semi-honest setting. For the sake of comparison, we omit the Big $O$ notations and simplify the complexity by retaining only the dominant terms. We use to denote protocols in the standard semi-honest model and to denote protocols requiring non-collusion assumption. pub: public-key operations; sym: symmetric cryptographic operations. We ignore the offline phase cost in all SK-MPSU protocols and the symmetric-key operations in all PK-MPSU protocols. [KS05] and [Fri07] use Paillier while [VCE22], [GNT23] and our PK-MPSU use ElGamal. $n$ is the set size. $m$ is the number of participants. $\lambda$ and $\sigma$ are computational and statistical security parameter respectively. $T$ is the number of AND gate in a SKE decryption circuit in [LG23]. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. In the typical setting, $n\leq 2^{24}$, $m\leq 32$, $\lambda=128$, $\sigma=40$, $T\approx 600$, $l\leq 128$ , $\gamma\leq 64$.
@@ 2.1 LG Protocol Revisit
We start by recalling the high-level idea of LG protocol. For the sake of simplicity, we focus here on the case of three parties $P_{1},P_{2},P_{3}$, whose inputs are $X_{1},X_{2},X_{3}$ respectively. We designate $P_{1}$ as the leader. Since $P_{1}$ already holds $X_{1}$, it needs to obtain the set difference $Y_{1}=(X_{1}\cup X_{2}\cup X_{3})\setminus X_{1}$ from $P_{2}$ and $P_{3}$.
Roughly speaking, their protocol enables $P_{2}$ to somehow secret-share $Y_{2}=X_{2}\setminus X_{1}$ and $P_{3}$ to somehow secret-share $Y_{3}=X_{3}\setminus(X_{1}\cup X_{2})$ among all parties. Since $\{Y_{2},Y_{3}\}$ is a partition of $Y_{1}$, each party holds a share of $Y_{1}$ in the order of $Y_{2},Y_{3}$ eventually, and any two parties cannot collude to obtain information of the last one’s inputs.
A naive approach to reconstructing $Y_{1}$ to $P_{1}$ is to let $P_{2}$ and $P_{3}$ send their shares to $P_{1}$ straightforwardly. However, in this way, $P_{1}$ can determine the party that each obtained element $x\in Y_{1}$ belongs. Their solution is to let the parties invoke multi-party secret-shared shuffle to randomly permute and re-share all shares. Multi-party secret-shared shuffle guarantees that any two parties have no knowledge of the permutation and all shares are refreshed after the invocation, hence the adversary is unaware of the correspondence between shares and the individual difference sets $Y_{2},Y_{3}$. Then $P_{2}$ and $P_{3}$ can send their shuffled shares to $P_{1}$, who reconstructs $Y_{1}$ and obtains the union by appending elements of $X_{1}$.
@@
LG protocol utilizes two main ingredients: (1) The secret-shared private membership test (ssPMT) [CO18, ZMS+21], where the sender $\mathcal{S}$ inputs a set $X$, and the receiver $\mathcal{R}$ inputs an element $y$. If $y\in X$, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of 1, otherwise they receive secret shares of 0. Liu and Gao proposed a multi-query ssPMT (mq-ssPMT), which supports the receiver querying multiple elements’ memberships of the sender’s set simultaneously. Namely, $\mathcal{S}$ inputs $X$, and $\mathcal{R}$ inputs $y_{1},\cdots,y_{n}$. $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $n$, where if $y_{i}\in X$, the $i$th bit is 1, otherwise 0. (2) A two-choice-bit version of random oblivious transfer (ROT), where the sender $\mathcal{S}$ and the receiver $\mathcal{R}$ each holds a choice bit $e_{0},e_{1}$. $\mathcal{S}$ receives two random messages $r_{0},r_{1}$. If $e_{0}\oplus e_{1}=0$, $\mathcal{R}$ receives $r_{0}$, otherwise $r_{1}$ [^2] This special ROT is identical to the standard 1-out-of-2 ROT, where $e_{0}$ is determined by $\mathcal{S}$ to indicate whether it would swap the order of $r_{0}$ and $r_{1}$. We use the standard ROT to formally describe our protocol specifications in the later sections. Here, we boil down a number of steps to the two-choice-bit ROT for facilitating illustration.. The following is to elaborate that how to utilize these ingredients to realize the secret-sharing processes.
@@
The process for $P_{2}$ to secret-share $Y_{2}$ is as follows: $P_{2}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$. For each item $x\in X_{2}$, $P_{2}$ and $P_{1}$ receive shares $e_{2,1}$ and $e_{1,2}$. If $x\in X_{1}$, $e_{2,1}\oplus e_{1,2}=1$, otherwise $e_{2,1}\oplus e_{1,2}=0$. Then $P_{2}$ acts as $\mathcal{S}$ and executes the two-choice-bit ROT with $P_{1}$. $P_{2}$ and $P_{1}$ each inputs $e_{2,1},e_{1,2}$. $P_{2}$ receives $r_{2,1}^{0},r_{2,1}^{1}$. If $e_{2,1}\oplus e_{1,2}=0$, $P_{1}$ receives $r_{2,1}^{0}$, otherwise $P_{1}$ receives $r_{2,1}^{1}$. $P_{1}$ sets the output as its share $s_{2,1}$. $P_{2}$ sets its share $s_{2,2}$ to be $r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ [^3] Suppose a lack of specific structure for set elements, then $P_{1}$ cannot distinguish a set element $x$ and a random value $r$ when it reconstructs the difference set. To address this, all parties append the hash value when secret sharing an element, i.e., sharing $x\|\mathsf{H}(x)$. It is provable that if the output length of $\mathsf{H}$ is sufficiently long (according to [LG23], at least $\sigma+\log(m-1)+\log n$ bits), the probability of existing $x$ that satisfies $x\|\mathsf{H}(x)=r$ is negligible. Therefore, $P_{1}$ can distinguish set elements from random values with overwhelming probability.. $P_{3}$ sets $s_{2,3}$ to 0. If $x\notin X_{1}$, $e_{2,1}\oplus e_{1,2}=0$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=r_{2,1}^{0}\oplus(r_{2,1}^{0}\oplus x\|\mathsf{H}(x))=x\|\mathsf{H}(x)$. Otherwise, $e_{2,1}\oplus e_{1,2}=1$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=r_{2,1}^{1}\oplus r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ is uniformly random. That is, $Y_{2}$ is secret-shared among all parties [^4] Since $X_{2}\setminus X_{1}$ merely contains the information of $X_{1}$ and $X_{2}$, it is sufficient to secret-share $X_{2}\setminus X_{1}$ among $P_{1}$ and $P_{2}$. Other difference sets are shared similarly., and the other elements in $X_{2}$ are masked by random values before being secret-share.
@@
We proceed to the process for $P_{3}$ to secret-share $Y_{3}$: $P_{3}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$ and $P_{2}$ separately. For each $x\in X_{3}$, $P_{3}$ and $P_{1}$ receive shares $e_{3,1}$ and $e_{1,3}$, while $P_{3}$ and $P_{2}$ receive shares $e_{3,2}$ and $e_{2,3}$. Then $P_{3}$ acts as $\mathcal{S}$ and executes the two-choice-bit ROT with $P_{1}$ and $P_{2}$. In the ROT between $P_{3}$ and $P_{1}$, $P_{3}$ inputs $e_{3,1}$ and $P_{1}$ inputs $e_{1,3}$. $P_{3}$ receives $r_{3,1}^{0},r_{3,1}^{1}$. If $e_{3,1}\oplus e_{1,3}=0$, $P_{1}$ receives $r_{3,1}^{0}$, otherwise $P_{1}$ receives $r_{3,1}^{1}$. $P_{1}$ sets the output as its share $s_{3,1}$. In the ROT between $P_{3}$ and $P_{2}$, $P_{3}$ inputs $e_{3,2}$ and $P_{2}$ inputs $e_{2,3}$. $P_{3}$ receives $r_{3,2}^{0},r_{3,2}^{1}$. If $e_{3,2}\oplus e_{2,3}=0$, $P_{2}$ receives $r_{3,2}^{0}$, otherwise $P_{2}$ receives $r_{3,2}^{1}$. $P_{2}$ sets the output as its share $s_{3,2}$. $P_{3}$ sets its share $s_{3,3}$ to be $r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)$. If $x\notin X_{1}$ and $x\notin X_{2}$, $e_{3,1}\oplus e_{1,3}=0$ and $e_{3,2}\oplus e_{2,3}=0$, $s_{3,1}\oplus s_{3,2}\oplus s_{3,3}=r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus(r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x))=x\|\mathsf{H}(x)$. Otherwise, there is at least one random value $r$ ($r_{3,1}^{1}$ or $r_{3,2}^{1}$) cannot be canceled out from the summation.
@@
The LG protocol described above has two main drawbacks: First, as the heart of LG protocol, mq-ssPMT is given a heavy instantiation which is fed much computation task through expensive general MPC machinery and renders it the bottleneck of the entire protocol. Second, LG protocol fails to achieve security against arbitrary collusion. In the following two sections, we are devoted to improving the efficiency of mq-ssPMT and enhancing LG protocol into standard semi-honest security, respectively.

### v4

@@ Abstract
Multi-party private set union (MPSU) protocol enables $m$ $(m>2)$ parties, each holding a set, to collectively compute the union of their sets without revealing any additional information to other parties. There are two main categories of multi-party private set union (MPSU) protocols: The first category builds on public-key techniques, where existing works require a super-linear number of public-key operations, resulting in their poor practical efficiency. The second category builds on oblivious transfer and symmetric-key techniques. The only work in this category, proposed by Liu and Gao (ASIACRYPT 2023), features the best concrete performance among all existing protocols, but still has super-linear computation and communication. Moreover, it does not achieve the standard semi-honest security, as it inherently relies on a non-collusion assumption, which is unlikely to hold in practice.
There remain two significant open problems so far: no MPSU protocol achieves semi-honest security based on oblivious transfer and symmetric-key techniques, and no MPSU protocol achieves both linear computation and linear communication complexity. In this work, we resolve both of them.
  • We propose the first MPSU protocol based on oblivious transfer and symmetric-key techniques in the standard semi-honest model. This protocol is 3.910.0×3.9-10.0$\times$3.9 - 10.0 × faster than Liu and Gao in the LAN setting. Concretely, our protocol requires only $4.4$ seconds in online phase for 3 parties with sets of $2^{20}$ items each.
@@
We implement our protocols and conduct an extensive experiment to compare the performance of our protocols and the state-of-the-art. To the best of our knowledge, our code is the first correct and secure implementation of MPSU that reports on large-size experiments.
@@ 1 Introduction
Over the last decade, there has been growing interest in private set operation (PSO), which consists of private set intersection (PSI), private set union (PSU), and private computing on set intersection (PCSI), etc. Among these functionalities, PSI, especially two-party PSI [PSZ14, KKRT16, PRTY19, CM20, PRTY20, RS21, RR22], has made tremendous progress and become highly practical with extremely fast and cryptographically secure implementations. Meanwhile, multi-party PSI [KMP+17, NTY21, CDG+21, BNOP22] is also well-studied. In contrast, the advancement of PSU has been sluggish until recently, several works proposed efficient two-party PSU protocols [KRTW19, GMR+21, JSZ+22, ZCL+23, CZZ+24a, JSZG24]. However, multi-party PSU has still not been extensively studied. In this work, we focus on PSU in the multi-party setting.
@@
Multi-party private set union (MPSU) enables $m$ $(m>2)$ mutually untrusted parties, each holding a private set of elements, to compute the union of their sets without revealing any additional information. MPSU and its variants have numerous applications, such as information security risk assessment [LV04], IP blacklist and vulnerability data aggregation [HLS+16], joint graph computation [BS05], distributed network monitoring [KS05], building block for private DB supporting full join [KRTW19], private ID [GMR+21], etc.
According to the underlying techniques, existing MPSU protocols [^1] We only consider the special-purpose solutions for MPSU, excluding those based on circuit-based generic techniques of secure computation, due to their unacceptable performance. can be divided into two categories:
  • PK-MPSU: This category is primarily based on public-key techniques, and has been explored in a series of works [KS05, Fri07, VCE22, GNT23]. A common drawback of these works is that each party has to perform a substantial number of public-key operations, leading to unsatisfactory practical efficiency.
@@
  • SK-MPSU: This category is primarily based on symmetric-key techniques, and includes only one existing work [LG23] to date. This work exhibits much better performance than all prior works but fails to achieve standard semi-honest security due to its inherent reliance on a non-collusion assumption, assuming the party who obtains the union (we call it leader hereafter) not to collude with other parties.
@@
Both categories share one common limitation: neither of them includes a protocol achieving linear computation and communication complexity [^2] In the context of MPSU, linear complexity means that the complexity per party scales linearly with the total size of all parties’ sets. In this paper, we consider the balanced setting where each party holds sets of equal size, thus linear complexity denotes the complexity per party to scale linearly with both the number of parties $m$ and the set size $n$. Meanwhile, following current conventions, linear complexity only considers the online phase.. Motivated by the context, we raise the following two questions:
Can we construct an MPSU protocol based on oblivious transfer and symmetric-key operations, without any non-collusion assumptions? Can we construct an MPSU protocol with both linear computation and linear communication complexity?
@@ 1.1 Our Contribution
In this work, we answer the above two questions affirmatively. Our contributions are summarized as follows:
  • Efficient batch ssPMT. We present a new primitive called batch secret-shared private membership test (batch ssPMT). Compared to multi-query secret-shared private membership test (mq-ssPMT), which is the technical core of the state-of-the-art MPSU protocol [LG23] (hereafter referred to as LG), batch ssPMT admits a much more efficient construction. Combined with hashing to bins, batch ssPMT can replace mq-ssPMT in the context of MPSU. Looking ahead, our batch ssPMT serves as a core building block in our two MPSU protocols, and significantly contributes to our speedup to LG.
  • SK-MPSU in standard semi-honest model. We generalize random oblivious transfer (ROT) into multi-party setting and present a new primitive called multi-party secret-shared random oblivious transfer (mss-ROT). Based on batch ssPMT and mss-ROT, we propose the first SK-MPSU in the standard semi-honest model. In addition to enhanced security, our SK-MPSU has superior online / total performance with a 3.910.0×3.9-10.0$\times$3.9 - 10.0 × / 1.27.8×1.2-7.8$\times$1.2 - 7.8 × improvement in the LAN setting.
  • PK-MPSU with linear complexity. Based on batch ssPMT and multi-key rerandomizable public-key encryption (MKR-PKE) [GNT23], we propose the first MPSU protocol with both linear computation and communication. Our PK-MPSU has the lowest overall communication costs with a factor of 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement compared to LG. Along the way, we find that the MPSU protocol of Gao et al. [GNT23] is insecure against arbitrary collusion and give a practical attack to demonstrate that it necessitates non-collusion assumption as well.  [^3] After we pointed out the security flaw of the GNT protocol [GNT23] with a concrete attack, the authors of [GNT23] contacted us and confirmed our attack. Subsequently, they updated their paper and revised their original protocol (c.f. Appendix B in the new version of [GNT23]) to a new one, which is similar to our PK-MPSU. The main difference lies in that they implement our batch ssPMT by invoking multiple instances of ssPMT separately, which renders their new protocol still has superlinear complexities. See also the summary of the relationship in their updated paper.
@@
  • 1.2 Related Works
    We review the existing semi-honest MPSU protocols below.
    • PK-MPSU. Kisser and Song [KS05] introduced the first MPSU protocol, based on polynomial representations and additively homomorphic encryption (AHE). This protocol requires a substantial number of AHE operations and high-degree polynomial calculations, so it is completely impractical.
@@
      Vos et al. [VCE22] proposed an MPSU protocol based on the bit-vector representations, where the parties collectively compute the union by performing the private OR operations on the bit-vectors, using ElGamal encryption. It shows poor concrete efficiency reported by [LG23] and the leader requires quadratic computation and communication in the number of parties.
      Recently, Gao et al. [GNT23] proposed an MPSU protocol, representing the most advanced MPSU in terms of asymptotic complexity. Unfortunately, their protocol turns out to be insecure against arbitrary collusion. We propose a practical attack to show that it requires the same non-collusion assumption as LG (see Appendix 0.A for details).
    • SK-MPSU. Recently, Liu and Gao [LG23] proposed a practical MPSU protocol based on oblivious transfer and symmetric-key operations. This protocol is several orders of magnitude faster than the prior works. For instance, when computing on datasets of $2^{10}$ element, it is 109×109$\times$109 × faster than [VCE22]. However, their protocol is not secure in the standard semi-honest model.
    • Other Related Works. Blanton et al. [BA12] proposed an MPSU protocol based on oblivious sorting and generic multi-party computation (MPC). The heavy dependency on general MPC leads to inefficiency.
@@
    Table 1: Asymptotic communication (bits) and computation costs of MPSU protocols in the semi-honest setting. For the sake of comparison, we omit the Big $O$ notations and simplify the complexity by retaining only the dominant terms. We use to denote protocols in the standard semi-honest model and to denote protocols requiring non-collusion assumption. pub: public-key operations; sym: symmetric cryptographic operations. $n$ is the set size. $m$ is the number of parties. $\lambda$ and $\sigma$ are computational and statistical security parameter. $T$ is the number of AND gate in a SKE decryption circuit in [LG23]. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. In the typical setting, $n\leq 2^{24}$, $m\leq 32$, $\lambda=128$, $\sigma=40$, $T\approx 600$, $l\leq 128$, $\gamma\leq 64$.
@@ 2.1 LG Revisit
    We start by abstracting the high-level idea of LG. For the sake of simplicity, we focus here on the case of three parties $P_{1},P_{2},P_{3}$, with respective inputs $X_{1},X_{2},X_{3}$. We designate $P_{1}$ as the leader. Since $P_{1}$ already holds $X_{1}$, it needs to obtain the set difference $Y_{1}=(X_{1}\cup X_{2}\cup X_{3})\setminus X_{1}$ from $P_{2}$ and $P_{3}$.
    LG can be summarized as a secret-sharing based MPSU framework with two phases: The first phase involves two secret-sharing processes, where $P_{2}$ somehow secret-shares $Y_{2}=X_{2}\setminus X_{1}$ and $P_{3}$ somehow secret-shares $Y_{3}=X_{3}\setminus(X_{1}\cup X_{2})$ among all parties. Since $\{Y_{2},Y_{3}\}$ is a partition of $Y_{1}$, as a result, the parties secret-share $Y_{1}$ in the order of $Y_{2},Y_{3}$. The second phase is to reconstruct $Y_{1}$ to $P_{1}$. In this case, a straightforward approach is insufficient because $P_{1}$ can identify which party each element $x\in Y_{1}$ originates from. The solution is to let the parties invoke multi-party secret-shared shuffle to randomly permute and re-share all shares. This shuffle ensures that any two parties have no knowledge of the permutation, thereby concealing the correspondence between the secret shares and individual difference sets $Y_{2},Y_{3}$. Afterwards, $P_{2}$ and $P_{3}$ can send their shuffled shares to $P_{1}$, who then reconstructs $Y_{1}$.
@@
    LG utilizes two ingredients to realize the secret-sharing processes of the framework: (1) The secret-shared private membership test (ssPMT) [CO18, ZMS+21], where the sender $\mathcal{S}$ inputs a set $X$, and the receiver $\mathcal{R}$ inputs an element $y$. If $y\in X$, $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of 1, otherwise secret shares of 0. Liu and Gao proposed a multi-query ssPMT (mq-ssPMT), which supports the receiver querying multiple elements’ memberships of the sender’s set simultaneously. Namely, $\mathcal{S}$ inputs $X$, and $\mathcal{R}$ inputs $y_{1},\cdots,y_{n}$. $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $n$, where if $y_{i}\in X$, the $i$th bit is 1, otherwise 0. (2) A two-choice-bit version of random oblivious transfer (ROT), where the sender $\mathcal{S}$ and the receiver $\mathcal{R}$ each holds a choice bit $e_{0},e_{1}$. $\mathcal{S}$ receives two random messages $r_{0},r_{1}$. If $e_{0}\oplus e_{1}=0$, $\mathcal{R}$ receives $r_{0}$, otherwise $r_{1}$ [^4] This special ROT is identical to the standard 1-out-of-2 ROT, where $e_{0}$ is determined by $\mathcal{S}$ to indicate whether to swap the order of $r_{0}$ and $r_{1}$.. The following is to elaborate the concrete constructions of the two processes.
@@
    The process for $P_{2}$ to secret-share $Y_{2}$: $P_{2}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$. For each item $x\in X_{2}$, $P_{2}$ and $P_{1}$ receive shares $e_{2,1}$ and $e_{1,2}$. If $x\in X_{1}$, $e_{2,1}\oplus e_{1,2}=1$, otherwise $e_{2,1}\oplus e_{1,2}=0$. Then $P_{2}$ acts as $\mathcal{S}$ and executes the two-choice-bit ROT with $P_{1}$. $P_{2}$ and $P_{1}$ each inputs $e_{2,1},e_{1,2}$. $P_{2}$ receives $r_{2,1}^{0},r_{2,1}^{1}$. If $e_{2,1}\oplus e_{1,2}=0$, $P_{1}$ receives $r_{2,1}^{0}$, otherwise $P_{1}$ receives $r_{2,1}^{1}$. $P_{1}$ sets the output as its share $s_{2,1}$. $P_{2}$ sets its share $s_{2,2}$ to be $r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ [^5] Suppose a lack of specific structure for set elements, then $P_{1}$ cannot distinguish a set element $x$ and a random value $r$ in the reconstruction. To address this, the parties append the hash value when secret sharing an element. . $P_{3}$ sets $s_{2,3}$ to 0. If $x\notin X_{1}$, $e_{2,1}\oplus e_{1,2}=0$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=x\|\mathsf{H}(x)$. Otherwise, $e_{2,1}\oplus e_{1,2}=1$, $s_{2,1}\oplus s_{2,2}\oplus s_{2,3}=r_{2,1}^{1}\oplus r_{2,1}^{0}\oplus x\|\mathsf{H}(x)$ is uniformly random. That is, $Y_{2}$ is secret-shared among all parties [^6] After the re-sharing of multi-party secret-shared shuffle., and the elements in $X_{1}\cap X_{2}$ are masked by random values.
@@
    The process for $P_{3}$ to secret-share $Y_{3}$: $P_{3}$ acts as $\mathcal{R}$ and executes the mq-ssPMT with $P_{1}$ and $P_{2}$ separately. For each $x\in X_{3}$, $P_{3}$ and $P_{1}$ receive shares $e_{3,1}$ and $e_{1,3}$, while $P_{3}$ and $P_{2}$ receive shares $e_{3,2}$ and $e_{2,3}$. Then $P_{3}$ acts as $\mathcal{S}$ and executes the two-choice-bit ROT with $P_{1}$ and $P_{2}$. In the ROT between $P_{3}$ and $P_{1}$, $P_{3}$ inputs $e_{3,1}$ and $P_{1}$ inputs $e_{1,3}$. $P_{3}$ receives $r_{3,1}^{0},r_{3,1}^{1}$. If $e_{3,1}\oplus e_{1,3}=0$, $P_{1}$ receives $r_{3,1}^{0}$, otherwise $P_{1}$ receives $r_{3,1}^{1}$. $P_{1}$ sets the output as its share $s_{3,1}$. In the ROT between $P_{3}$ and $P_{2}$, $P_{3}$ inputs $e_{3,2}$ and $P_{2}$ inputs $e_{2,3}$. $P_{3}$ receives $r_{3,2}^{0},r_{3,2}^{1}$. If $e_{3,2}\oplus e_{2,3}=0$, $P_{2}$ receives $r_{3,2}^{0}$, otherwise $P_{2}$ receives $r_{3,2}^{1}$. $P_{2}$ sets the output as its share $s_{3,2}$. $P_{3}$ sets its share $s_{3,3}$ to be $r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)$. If $x\notin X_{1}$ and $x\notin X_{2}$, $e_{3,1}\oplus e_{1,3}=0$ and $e_{3,2}\oplus e_{2,3}=0$, $s_{3,1}\oplus s_{3,2}\oplus s_{3,3}=x\|\mathsf{H}(x)$. Otherwise, there is at least one random value ($r_{3,1}^{1}$ or $r_{3,2}^{1}$) cannot be canceled out from the summation that makes it random.
@@
    The above outlines how LG works. The protocol has two main drawbacks: First, the core component, mq-ssPMT heavily relies on expensive general MPC machinery, which becomes the bottleneck of the entire protocol. Second, LG fails to achieve security against arbitrary collusion. The next two sections are devoted to addressing these two drawbacks.